OpenStudy (anonymous):

how do you get the value of sin35(degrees) MANUALLY?

OpenStudy (amistre64):

by luckily finding 2 angles that you know of that will add or subtract to it

OpenStudy (lgbasallote):

sin(60-45)

OpenStudy (amistre64):

sin(a+b) = sin(a)cos(b)+sin(a)cos(b)

OpenStudy (amistre64):

thing is 60-45 not= 35 :)

OpenStudy (anonymous):

I think that is is equal to sin(20 +15) an I know how to solve for sin 15.but how do I get the value of sin20? O_O

OpenStudy (anonymous):

sin(60-45) =sin60cos45+sin45cos60

OpenStudy (anonymous):

60-45 is never 35 :)

OpenStudy (lgbasallote):

oh yes.. my mistake :) maybe sin(30 + 30/6)? = sin30cos30/6 + cos30sin30/6 but it involves a lot of theorems

OpenStudy (phi):

In the old days, you would get out a quill pen and start figuring... with x in radians, this has a relative error of less than 1.9% \[ sin(x) =\frac{16x(\pi-x)}{5\pi^{2}-4x(\pi-x)} \] See http://en.wikipedia.org/wiki/History_of_trigonometry Indian Mathematics Or use Newton's series definition (again, x in radians) \[ sin(x)= x - \frac{x^{3}}{3!} + \frac{x^{5}}{5!} - \frac{x^{7}}{7!} + …\] Test: sin(35)= 0.573576436 by a Casio calculator with pi= 3.141592654 x= 35*pi/180= 0.610865238 radians (tedious if done by hand, but doable) Using formula (1) 16*0.610865238*(pi-0.610865238)/(5*pi*pi-4*0.610865238*(pi-0.610865238) = 0.573041637 Using 4 terms of Newton's series 0.610865238-0.610865238^3/6 + 0.610865238^5/120-0.610865238^7/5040 =0.573576404 Using 5 terms: 0.573576404+0.610865238^9/362880 =0.573576437